The sum of two numbers is 1. Their difference is 3. What is the smaller of the two numbers?: The sum of two numbers is 1. Their difference is 3. What is the smaller of the two numbers?
Answer by jim_thompson5910(8211)

"The sum of two numbers is 1." translates to x+y=1


"Their difference is 3." translates to x-y=3



Start we have the system of equations:

system(x+y=1,x-y=3)


Let's solve the system by use of substitution


Now in order to solve this system by using substitution, we need to solve (or isolate) one variable. I'm going to solve for y.




So let's isolate y in the first equation

x+y=1 Start with the first equation


y=1-x Subtract x from both sides


y=-x+1 Rearrange the equation



---------------------

Since y=-x+1, we can now replace each y in the second equation with -x+1 to solve for x



x-highlight((-x+1))=3 Plug in y=-x+1 into the first equation. In other words, replace each y with -x+1. Notice we've eliminated the y variables. So we now have a simple equation with one unknown.



x+x-1=3 Distribute the negative


2x-1=3 Combine like terms on the left side


2x=3+1Add 1 to both sides


2x=4 Combine like terms on the right side


x=(4)/(2) Divide both sides by 2 to isolate x



x=2 Divide





-----------------First Answer------------------------------


So the first part of our answer is: x=2









Since we know that x=2 we can plug it into the equation y=-x+1 (remember we previously solved for y in the first equation).



y=-x+1 Start with the equation where y was previously isolated.


y=-(2)+1 Plug in x=2


y=-2+1 Multiply


y=-1 Combine like terms



-----------------Second Answer------------------------------


So the second part of our answer is: y=-1









-----------------Summary------------------------------

So our answers are:

x=2 and y=-1



So the smaller of the two numbers is -1
Question 145788: The sum of two numbers is 1. Their difference is 3. What is the smaller of the two numbers?: The sum of two numbers is 1. Their difference is 3. What is the smaller of the two numbers?
Answer by jojo14344(326)

Well, Let "x" be the 1st #
"y" be the 2nd number
In 1st condition, 'x+y=1' -------------eqn 1
In 2nd condition, 'x-y=3' -------------eqn 2
In eqn 1 we get 'x= 1-y' right? substitute this to eqn 2,
1-y-y=3
2y=-2
y= -1
And the other # as per eqn 1, x+(-1)=1
x=1+1 = 2
So, the smaller # is y= -1
In doubt? Go either eqn, okay we use eqn 2,
x-y=3
2-(-1)=3
2+1=3
3=3
Thank you,
Jojo